OpenStudy (anonymous):

Is the simplified form of 2square root of 3 − 2square root of 3 rational? Yes No

OpenStudy (anonymous):

i think it is because 2^3-2^3= 8-8=0

OpenStudy (solomonzelman):

well, actually, \(\large\color{#000000 }{ \displaystyle 2\sqrt{3} -\displaystyle 2\sqrt{3}}\) , but the result is the same.